A 5-kg dog was infused with a drug with the strength of 20 mg/ml. In the last 12 hours the dog has received 15 ml. Calculate the rate of infusion. .
Answer #1

A 5-kg dog was infused with a drug with the strength of 20 mg/ml. in the last 12 hours the dog has received 15 ml. calculate the rate of infusion.

Solution:- Given that,

C.F.= (15/12) = 1.25

dog weight = 5-kg ; strength = 20 mg/ml,

the last 12 hours the dog has received 15 ml

the rate of infusion =?

5 kg * 20 mg/ml = 100 ml

100 ml* (15) = 1500 ml/hr

1500 ml/hr *1.25 = 1875 ml/ hr

the rate of infusion = 1875 ml/ hr